# Perverse microsheaves

### v2

@@ 5. Microsheaves on complex contact manifolds
where $\tilde{V}$ is an exact complex symplectic manifold with $1$-form $\lambda_{\tilde{V}}$, and $\tilde{V}/\mathbb{R}_{+}$ carries a $\mathbb{C}^{*}$ of real contact forms $\xi_{\hbar}:=\operatorname{re}(\hbar\lambda_{\tilde{V}})$. Evidently $(\tilde{V},\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ is the real symplectization of $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. For specificity we fix $\hbar=0$ (but see Remark 5.7).
For a complex Legendrian $L\subset V$, we similarly write $\tilde{L}:=\pi^{-1}(L)$ and $\tilde{L}/\mathbb{R}_{+}:=p^{-1}(L)$.
@@
Let $V$ be a complex contact manifold and $\tilde{V}$ its symplectization. Per Lemma 3.2 and Remark 3.4, there is a canonical grading for $\tilde{V}$. Choose $C$-orientation data $o$ on $V$ and pull it back to $\tilde{V}$; combining this with with the canonical grading, we obtain Maslov data. We also write $o$ for this Maslov data, and correspondingly $\mu sh_{\tilde{V},o}$ for the corresponding sheaf of categories on $\tilde{V}$. In case $C=R-mod$ for a commutative ring $R$, then we have available the canonical orientation data; when we make this choice, we omit $o$ from the notation entirely, writing simply $\mu sh_{\tilde{V}}$.
For $\tilde{\Lambda}\subset\tilde{V}$ a $\mathbb{R}_{>0}$-invariant complex-analytic subset which is the closure of its smooth Lagrangian locus, we write $\mu sh_{\tilde{\Lambda}}\subset\mu sh_{\tilde{V}}$ for the subsheaf of full subcategories on objects supported in $\tilde{\Lambda}$. We write $\mu sh_{\tilde{V},\mathbb{C}-c,o}$ for the union of all such $\mu sh_{\tilde{\Lambda},o}$.
@@
For any fixed subanalytic Lagrangian $\Lambda$, the category $sh_{\Lambda}(M)$ is presentable, but $sh_{\mathbb{C}-c}(M)$ is not (arbitrary sums of constructible sheaves certainly need not be constructible). The situation for the microsheaf categories is entirely analogous. The distinction is rarely relevant: for example, in the present article, while we state theorems for $\mu sh_{\tilde{V},\mathbb{C}-c,o}$, their proofs quickly reduce to statements about $\mu sh_{\tilde{\Lambda},o}$.
@@
A closed and complex analytic subset $\tilde{\Lambda}\subset\tilde{V}$ is $\mathbb{R}_{>0}$-invariant iff it is $\mathbb{C}^{*}$-invariant. We use the formulation above because we may want to consider open sets which are $\mathbb{R}_{>0}$-invariant but not $\mathbb{C}^{*}$-invariant, and so need a corresponding local notion.
Remark 5.4.
Per Remark 4.17, we could equivalently have defined a category of microsheaves on the real contact manifold $\tilde{V}/\mathbb{R}_{>0}$ and pulled it back along $q:\tilde{V}\to\tilde{V}/\mathbb{R}_{+}$; our choice of formulation above is solely to avoid the minor cognitive dissonance of passing constantly to a non-complex manifold.
Remark 5.5.
One could also consider the subsheaf of full subcategories of $\mu sh_{\tilde{V}}$ on objects whose microsupport is complex-analytic, but this would e.g. not be closed under taking summands.
Lemma 5.6.
Let $f:U\to V$ be a complex contact embedding, and $\tilde{f}:\tilde{U}\to\tilde{V}$ its lift to the complex symplectization. Then there is a canonical isomorphism $\mu sh_{U,f^{*}o}=f^{*}\mu sh_{V,o}$.
In particular, when $o$ is the canonical orientation datum, there is a canonical isomorphism $\mu sh_{U}=f^{*}\mu sh_{V}$.
Proof.
It suffices to argue that the Maslov data of $V$ pulls back to the Maslov data of $U$, up to canonical homotopy. The orientation data is respected by definition, while the grading is canonically induced by the complex structure, hence behaves naturally under the above pullback. ∎
Remark 5.7.
$\tilde{V}$ is a $\mathbb{C}^{*}$-principal bundle over $V$, and the natural $\mathbb{C}^{*}$ action evidently satisfies $\hbar^{*}\lambda_{\tilde{V}}=\hbar^{-1}\lambda_{\tilde{V}}$ for $\hbar\in\mathbb{C}^{*}$. In particular, multiplication by $\hbar$ descends to a real contactomorphism $(\tilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)\cong(\tilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)$. The action of $\hbar$ is also canonically trivial on Maslov data, as this is pulled back from $V$. Thus by similar reasoning as in Lemma 5.6, there is a canonical isomorphism $\hbar^{*}\mu sh_{(\tilde{V},\hbar\lambda)}\cong\mu sh_{(\tilde{V},\lambda)}$.
Lemma 5.8.
Let $M$ be a complex manifold. Consider the complex contact manifold $V=\mathbb{P}T^{*}M$, and the corresponding $\tilde{V}=T^{\circ}M$, with canonical orientation data. A spin structure on $M$ determines an equivalence of sheaves of categories on $T^{\circ}M$: $\mu sh\cong\mu sh_{T^{\circ}M}$ where the left hand side is as in (9) and the right hand side is defined as in Definition 5.1.
Proof.
Follows from Corollary 4.13 by Lemma 3.11 and Lemma 3.12. ∎
Even if we had used some arbitrary choice of Maslov data $\mu$ on $V$, not necessarily coming from orientation data and the canonical complex grading, the (complex) Darboux theorem and general functoriality of microsheaves would nevertheless imply that $\mathbb{P}\mu sh_{V,\mu}$ is locally isomorphic to any $\mathbb{P}\mu sh_{\mathbb{P}T^{*}M}$ of the correct dimension, since the space of Maslov data, being a torsor for $Maps(V,BPic(C))$, is certainly connected when $V$ is contractible. However, the resulting isomorphism is not canonical, even locally. What we gain from using the canonical grading is:
Lemma 5.9.
Let $U$ be a contractible complex contact manifold, and $u:U\to V$ be any complex contact embedding. Then for any choices of $C$-orientation data $o,o^{\prime}$, there is an isomorphism $\mu sh_{U,o}\cong u^{*}\mu sh_{V,o^{\prime}}$ which is canonical up to a non-canonical natural transformation.
Proof.
Using Lemma 5.6, a homotopy of orientation data $u^{*}o^{\prime}\sim o$ will determine such an isomorphism. This is defined up to the action of loops in the space of orientation data; as $U$ is contractible, this is just (by definition of $C$-orientation data, see Remark 4.20) the connected component of the identity $Pic(C)_{0}\subset Pic(C)$, which acts by natural transformations. ∎
If $\Lambda\subset V$ is a Legendrian in a real contact manifold equipped with appropriate orientation and grading data, it follows, ultimately from Proposition 4.6, that there is a non-canonical isomorphism $(\mu sh_{\Lambda})_{p}\xrightarrow{\sim}C$ at any smooth point $p\in\Lambda$. Any choice of such an isomorphism is termed a microstalk functor; there is a $Pic(C)$ worth of choices.
In (10), we saw how to define a distinguished microstalk functor for Legendrians which occur as conormals to embedded submanifolds, where the orientation data is induced by the fiber polarization. However, leaving this class of submanifolds, the torsor of microstalk functors has some monodromy (classified by the map $\mu_{C}$ from Theorem 4.14), typically involving nontrivial shifts.
A central observation of this paper is that the situation is better in the complex world:
Definition-Proposition 5.10.
Let $V$ be a complex contact manifold equipped with some orientation data $o$. Let $L\subset V$ be a complex Legendrian, and $p\in L$ a smooth point with lift $\tilde{p}\in\tilde{L}$.
Choose a codimension zero embedding $\iota:Nbd(p)\hookrightarrow\mathbb{P}T^{*}M$ for a complex manifold $M$, such that the projection $L\cap Nbd(p)\to M$ is an embedding onto its image. Consider the composition of isomorphisms:
$C\xrightarrow{\omega_{\iota(\tilde{p})}}(\mu sh_{\iota(\tilde{L})})_{\iota(\tilde{p})}=(\mu sh_{\iota(\tilde{L}),\phi})_{\tilde{p}}\cong(\mu sh_{\tilde{L},o})_{\tilde{p}}$
here, the first arrow is (10), the second is (the $T^{\circ}M$ version of) Corollary 4.13, and the last is Lemma 5.9. We have seen the first two maps are independent of choices, and the final map depends on choices only through (noncanonical) natural transformation. We denote the composition by $\omega_{\tilde{p}}$, and, henceforth, when we refer in this context to microstalk functors, we mean only those of the form $\omega_{\tilde{p}}^{-1}$.
In particular, given $F\in(\mu sh_{\tilde{L},o})_{\tilde{p}}$, the object $\omega_{\tilde{p}}^{-1}(F)\in C$ is well defined (i.e. independent of the embedding $\iota$). We term this object the microstalk at $\tilde{p}$.
Definition 5.11.
Fix a subcategory $D$ of our coefficient category $C$, and fix $C$-orientation data $o$ (or more generally, $(C,D)$-orientation data). Let $V$ be a complex contact manifold. We write $(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{D}\subset\mu sh_{\tilde{V},\mathbb{C}-c,o}$ for the full subcategory on objects $F$ such that at every smooth point $p\in supp(F)$, one has $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})\in D$.
Remark 5.12.
By the same argument as in Lemma 5.9, the isomorphism class of $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})$ is constant in the smooth locus of the microsupport. So we need only check $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})\in D$ at one point in each connected component of the smooth locus.
Legendrians give objects:
Proposition 5.13.
Let $V$ be a complex contact manifold and $L\subset V$ a complex Legendrian. Then relative orientation data on $L$ (e.g. a spin structure if we use the canonical orientation datum on $V$) determines an equivalence of categories $\mu sh_{\tilde{V},\mathbb{C}}|_{\tilde{L}}\cong Loc_{\tilde{L}}$ where $Loc_{\tilde{L}}$ is the sheaf of categories of local systems.
Proof.
As described in Remark 4.16, it suffices to provide secondary Maslov data. The grading part comes from Lemma 3.11. Lemma 3.12 explains why, for the canonical orientation datum, the corresponding notion of relative orientation data reduces is identified with spin structure. ∎
We now introduce sheaves of categories which will carry the microlocal perverse $t$-structure.
Definition 5.14.
Let $V$ be a complex-contact manifold. Given an object $K\in\pi_{*}\mu sh_{\tilde{V}}(U),U\subset V,$ and a point $p\in ss(K)\cap U$, we say that $K$ is representable at $p$ by a sheaf if there exists a contact embedding $\iota:Nbd(p)\hookrightarrow\mathbb{P}^{*}M$ such that (the image of) $K$ in $(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$ is represented by a sheaf $\mathcal{F}_{K}\in sh(M)$.222i.e. the image of $K$ in $(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$ coincides with the image of $\mathcal{F}_{K}$ under the quotient $sh(M)\to(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$. Here $M$ is any complex manifold and we recall that $\pi$ is the projection $T^{*}M-0_{M}=\widetilde{\mathbb{P}^{*}M}\to\mathbb{P}^{*}M$ – not projection to the base. We say that $K$ is pointwise representable by a sheaf if it is so at all points in its microsupport.
Remark 5.15.
By the discussion around [15, (11.4.8)] this definition is independent of the choice of the Darboux chart.
Definition 5.16.
Let $V$ be a contact manifold, $o$ orientation data, and $\Lambda$ a complex (singular) Legendrian. We denote by:
$\mathbb{P}\mu sh_{V,o}\subset\pi_{*}\mu sh_{\tilde{V},o}\qquad\qquad\mathbb{P}\mu sh_{\Lambda,o}\subset\pi_{*}\mu sh_{\pi^{-1}(\Lambda),o}\qquad\qquad\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}\subset\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c,o}$
the sheaves of full subcategories of $\pi_{*}\mu sh_{\tilde{V},o}$ (resp. $\pi_{*}\mu sh_{\pi^{-1}(\Lambda),o}$, $\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c,o}$) on objects which are pointwise representable by sheaves.
Remark 5.17.
It is a tautology that $\mathbb{P}\mu sh_{V,\mathbb{C}-c}$ is a union over all complex (singular) legendrians $\Lambda$ of $\mathbb{P}\mu sh_{\Lambda}$.
Remark 5.18.
We could equivalently have pushed forward the corresponding sheaves of categories defined on the circle bundle $\tilde{V}/\mathbb{R}_{>0}$.
6. Microsheaves on complex exact symplectic manifolds
Let $X=(X,\lambda)$ be an exact complex symplectic manifold. Consider the contact thickening $(X\times\mathbb{C},\lambda+dz)$ and the associated symplectization
(13) $(X\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz)).$
Let $V_{\mathbb{R}}$ denote the Liouville vector field for $(X,\lambda)$ and let $I$ denote the almost-complex structure induced by complex multiplication. Integrating the flow of $Z$ in the first component and the flow of $V_{S^{1}}:=IV_{\mathbb{R}}$ in second, we obtain an action $A:\mathbb{R}\times\mathbb{R}\times X\to X$ which satisfies $A_{(t,\theta)}^{*}\lambda=e^{t+2\pi i\theta}\lambda$.
If $U\subset\mathbb{C}^{*}$ is contractible, then we can define a action $U\times X\to X$ by lifting $U$ to $\mathbb{R}\times\mathbb{R}$ by the covering map $(t,\theta)\mapsto e^{t+2\pi i\theta}$. In many situations, the flow in the $IZ$ direction factors through $k\mathbb{Z}\subset\mathbb{R},k\geq 1$; in this case, we can lift the action via the covering map $(t,\theta)\mapsto e^{k(t+2\pi i\theta)}$ to define a weight $k$ action of $\mathbb{C}^{*}$.
(14) $(X\times\mathbb{C}\times U,\lambda+wdz)$ $\xrightarrow{\simeq}(X\times\mathbb{C}\times U,w(\lambda+dz))$
$(x,z,w)$ $\mapsto(A_{w^{-1}}(x),z,w)$
We let $\pi:X\times\mathbb{C}\times\mathbb{C}^{*}\to X\times\mathbb{C}$ be the projection which forgets the last component.
Recall that in 5.10, we show that, for a complex contact manifold $V$ and a complex Legendrian $L\subseteq V$, we can fix a canonical identification $(\mu sh_{\tilde{L}})_{\tilde{p}}\simeq C$ where $\tilde{L}$ is its Lagrangian lift and $\tilde{p}$ a smooth point. The above change of coordinates implies that a similar choice can be made for the symplectic case.
Lemma 6.1.
Suppose now that $L\subset X$ is a (possibly singular) conic complex-Lagrangian. Given $p\in L$, we have
$(\mu sh_{L})_{p}=\left(\mu sh_{\pi^{-1}(L\times\{0\})}\right)_{(p,0,1)}$
where the left hand side is defined as in [22, Section 9.1], for any exact symplectic manifolds, and the right hand side is defined in Definition 5.1, only for symplectizations.
Proof.
We have $\pi^{-1}(L\times\{0\})=L\times T^{\circ}_{0}\mathbb{C}=L\times\{0\}\times\mathbb{C}^{*}$. As a consequence of the local splitting (14), we have
(15) $\left(\mu sh_{\pi^{-1}(L\times\{0\})}\right)_{(p,0,1)}\simeq(\mu sh_{L})_{p}\otimes(\mu sh_{T^{\circ}_{0}\mathbb{C}})_{(0,1)}\simeq(\mu sh_{L})_{p}$
We now define the microstalk functor for exact complex-symplectic manifolds.
Definition 6.2.
Let $(X,\lambda)$ be an exact complex-symplectic manifold equipped with some orientation data $o$. Let $L\subset X$ a (germ of a) conical complex Lagrangian submanifold. For $p\in L$ a smooth point, we have a microstalk funtor defined as the composition
$C\to\left(\mu sh_{\pi^{-1}(L\times\{0\}),o}\right)_{(p,0,1)}=(\mu sh_{L,o})_{p}$
where the first arrow is the microstalk functor introduced in 5.10 and the equality is given by (15) above.
Proof of Theorem 1.2.
Per Lemma 3.2 and Remark 3.4, there is a canonical grading for $\tilde{V}$. Choose the canonical orientation data. Combining these, we obtain Maslov canonical data and hence a canonical sheaf of categories $\mu sh_{X}$. Note that here we are viewing $X=(X,\operatorname{re}\lambda)$ as a real exact symplectic manifold, so that one can define $\mu sh_{X}$ as in [22, Section 9.1]. More precisely, we identify $X$ with the subset $X\times\{0\}$ of its contactization $X\times\mathbb{R}$, and we then define $\mu sh_{X}$ as the restriction $\mu sh_{X\times\{0\}}|_{X\times\{0\}}$.
Finally, if $L$ is a smooth smooth conic complex spin Lagrangian, the canonical equivalence $\mu sh_{W,L}(-)\simeq loc(L)$ is established exactly like Proposition 5.13. ∎
Given an exact complex-symplectic manifold $(X,\lambda)$, consider the contactization $(X\times\mathbb{C},\lambda+dz)$. This is a complex-contact manifold, and so we can consider $\mathbb{P}\mu sh_{X\times\{0\}}\subset\mathbb{P}\mu sh_{X\times\mathbb{C}}$, which is the full subcategory on objects microsupported on $X\times\{0\}$.
We now have two sheaves of categories on $X$, namely $\mu sh_{X}$ and $\mathbb{P}\mu sh_{X\times\{0\}}$. They are not the same: for example, if $X$ is a point, then $\mu sh_{X}=C$ while $\mathbb{P}\mu sh_{X\times\{0\}}$ is the category of local systems on $\mathbb{C}^{*}$. In general, we do not know how to relate $\mu sh_{X}$ and $\mathbb{P}\mu sh_{X\times\{0\}}$. However:
Theorem 6.3.
When the Liouville vector field of $(X,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, there is a natural $\mathbb{Z}=\Omega S^{1}$-linear structure on $\mathbb{P}\mu sh_{X\times\{0\}}$, and there are fully faithful embeddings
(16) $\mathbb{P}\mu sh_{X\times\{0\}}\hookrightarrow\mu sh_{X}\otimes\Omega S^{1}$
(17) $\mu_{\mathbb{C}}sh_{X}:=\mathbb{P}\mu sh_{X\times\{0\}}\otimes_{\Omega S^{1}}\bullet\hookrightarrow\mu sh_{X}$
Of course, when working over a discrete ring, we may replace $\Omega S^{1}$ and $\bullet$ by their linearizations $\mathbb{Z}[\mathbb{Z}]=\mathbb{Z}[t,t^{-1}]$ and $\mathbb{Z}$.
Proof.
The change of variables (14) is global:
(18) $(\widetilde{X\times\mathbb{C}}=X\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz))$ $\xrightarrow{\simeq}(X\times\mathbb{C}\times\mathbb{C}^{*},\lambda+wdz)$
$\mu sh_{X\times 0\times\mathbb{C}^{*}}=\mu sh_{X}\boxtimes\mu sh_{\mathbb{C}^{*}\subset T^{*}\mathbb{C}^{*}}.$
Pushing forward and composing with the embedding $\mathbb{P}\mu sh_{X\times 0}\hookrightarrow\pi_{*}\mu sh_{X\times 0\times\mathbb{C}^{*}}$ gives the first assertion, from which the second follows formally. ∎
Remark 6.4.
(17) amounts to taking invariants of a topological $\mathbb{C}^{*}$-action; see [20, Sec. 6]. With respect to the split symplectic form (i.e. the right hand side of (18)), this action is $\theta\cdot(x,z,w)\mapsto(x,\theta z,\theta^{-1}w)$. With respect to the coupled symplectic form (i.e. the left hand side of (18)), this corresponds to the diagonal $\mathbb{C}^{*}$ action $\theta(x,z,w)=(\theta\cdot x,\theta z,\theta^{-1}w)$. [^3] Note the slight abuse of notation: $\theta\cdot x$ refers to the image of $x\in X$ under the time $\theta$ flow of the Liouville vector field, which integrates by assumption to a $\mathbb{C}^{*}$ action. Observe that this is just the lift to the symplectization of the $\mathbb{C}^{*}$-action $\theta\cdot(x,z)=(\theta\cdot x,\theta z)$ on $X\times\mathbb{C}$ by contactomorphism.
Remark 6.5.
If $(X,\omega)$ admits a weight $\mathbb{C}^{*}$-action of weight $k>1$, we expect a related statement, except that $\mathbb{P}\mu sh_{X}$ should be a full subcategory of $(\mu sh_{X})^{\mathbb{Z}/k}$.

### v3

@@ 5.3. Microsheaves on complex contact manifolds and symplectic manifolds
where $\widetilde{V}$ is an exact complex symplectic manifold with holomorphic $1$-form $\lambda_{\widetilde{V}}$. For any subset (typically complex Legendrian) $L\subset V$, we similarly write $\widetilde{L}:=\pi^{-1}(L)$ and $\widetilde{L}/\mathbb{R}_{+}:=p^{-1}(L)$.
Letting $\hbar\in\mathbb{C}^{*}$ act by the $\mathbb{C}^{*}$ principal bundle structure of $\widetilde{V}$ over $V$, we have $\hbar^{*}\lambda_{\widetilde{V}}=\hbar\lambda_{\widetilde{V}}$ for $\hbar\in\mathbb{C}^{*}$. In particular, multiplication by $\hbar$ descends to a real contactomorphism $(\widetilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)\cong(\widetilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\hbar\lambda)$.
Per Definition 3.8, $(\widetilde{V},\operatorname{re}\lambda)$ carries canonical Maslov data $\eta_{can}$. We write
$\mu sh_{\widetilde{V}}\coloneqq\mu sh_{(\widetilde{V},\operatorname{re}\lambda),\eta_{can}}.$
Here we have, and will henceforth, set $\hbar=1$. This is no loss of generality:
Lemma 5.12.
There is a canonical isomorphism $\hbar^{*}\mu sh_{(\widetilde{V},\operatorname{re}\hbar\lambda)}\cong\mu sh_{(\widetilde{V},\operatorname{re}\lambda)}$.
Proof.
The canonical Maslov datum is pulled back from $V$, so the action of the contactomorphism given by multiplication by $\hbar$ is also canonically trivial on it. ∎
More generally, we can consider any Maslov data $\eta$ on $\widetilde{V}$ and the associated sheaf $\mu sh_{\widetilde{V},\eta}$.
Remark 5.13.
The sheaf of categories $\mu sh_{\widetilde{V}}$ is the pullback of the corresponding object on the real contact manifold $\widetilde{V}/\mathbb{R}_{>0}$; we adopt our present formulation solely to avoid the minor cognitive dissonance of passing constantly to a non-complex manifold.
A closed and complex analytic subset $\widetilde{\Lambda}\subset\widetilde{V}$ is $\mathbb{R}_{>0}$-invariant iff it is $\mathbb{C}^{*}$-invariant, and hence the preimage of some $\Lambda\subset V$. We will however also write $\widetilde{\Lambda}$ for complex and $\mathbb{R}_{>0}$-invariant (but not necessarily closed or $\mathbb{C}^{*}$-invariant) subsets of $\widetilde{V}$. We say such a subset is Lagrangian if it is contained in the closure of its smooth and Lagrangian locus.
@@
We write $\mu sh_{\widetilde{V},\mathbb{C}-c,\eta}\subset\mu sh_{\widetilde{V},\eta}$ for the sheaf of full subcategories on objects with complex Lagrangian microsupport.
By definition, for $\mathbb{R}_{>0}$-conic subsets $\Omega\subset\widetilde{V}$, the space of sections $\mu sh_{\widetilde{V},\mathbb{C}-c,o}$ is the union of the categories $\mu sh_{\widetilde{\Lambda},o}(\Omega)$, where $\widetilde{\Lambda}$ varies over $\mathbb{R}_{>0}$-conic complex Lagrangian subsets $\widetilde{\Lambda}$ which are closed in $\Omega$.
@@
For any fixed subanalytic Lagrangian $\Lambda$, the category $sh_{\Lambda}(M)$ is presentable, but $sh_{\mathbb{C}-c}(M)$ is not (arbitrary sums of constructible sheaves certainly need not be constructible). The situation for the microsheaf categories is entirely analogous. The distinction is rarely relevant: for example, in the present article, while we state theorems for $\mu sh_{\widetilde{V},\mathbb{C}-c,o}$, their proofs quickly reduce to statements about $\mu sh_{\widetilde{\Lambda},o}$.
We now discuss complex symplectic manifolds. Let $W=(W,\lambda)$ be an exact complex symplectic manifold, consider the contact thickening $(W\times\mathbb{C},\lambda+dz)$ and the associated symplectization
(20) $(W\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz)).$
Let $V_{\mathbb{R}}$ denote the Liouville vector field for $(W,\lambda)$ and let $I$ denote the almost-complex structure induced by complex multiplication. Integrating the flow of $V_{\mathbb{R}}$ in the first component and the flow of $V_{S^{1}}:=IV_{\mathbb{R}}$ in second, we obtain an action $A:\mathbb{R}\times\mathbb{R}\times W\to W$ which satisfies $A_{(t,\theta)}^{*}\lambda=e^{t+2\pi i\theta}\lambda$.
If $U\subset\mathbb{C}^{*}$ is contractible, then we can define an action $U\times W\to W$ by lifting $U$ to $\mathbb{R}\times\mathbb{R}$ by the covering map $(t,\theta)\mapsto e^{t+2\pi i\theta}$. In many situations, the flow in the $IZ$ direction factors through $k\mathbb{Z}\subset\mathbb{R},k\geq 1$; in this case, we can lift the action via the covering map $(t,\theta)\mapsto e^{k(t+2\pi i\theta)}$ to define a weight $k$ action of $\mathbb{C}^{*}$.
(21) $(W\times\mathbb{C}\times U,\lambda+wdz)$ $\xrightarrow{\simeq}(W\times\mathbb{C}\times U,w(\lambda+dz))$
$(x,z,w)$ $\mapsto(A_{w^{-1}}(x),z,w)$
Given an exact complex-symplectic manifold $(W,\lambda)$, consider the contactization $(W\times\mathbb{C},\lambda+dz)$. This is a complex-contact manifold, and so we can consider $\mu sh_{W\times\{0\}}\subset\mu sh_{W\times\mathbb{C}}$, which is the full subcategory on objects microsupported on $W\times\{0\}$. We also write $\mu sh_{W\times\{0\},\mathbb{C}-c}$ for the full subcategory of $\mu sh_{W\times\{0\}}$ on objects with complex constructible support.
We now have two sheaves of categories on $W$, namely $\mu sh_{W}\coloneqq\mu sh_{(W,\operatorname{re}\lambda)}$ from Definition 4.25 and $\mu sh_{W\times\{0\}}$. They are not the same: for example, if $W$ is a point, then $\mu sh_{W}=C$ while $\mu sh_{W\times\{0\}}$ is the category of local systems on $\mathbb{C}^{*}$.
Theorem 5.16.
Suppose that the Liouville vector field of $(W,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action. Let $\gamma_{\mathbb{C}}:W\to W$ be the set-theoretic identity, where the source is endowed with the Euclidean topology and the target with the $\mathbb{C}^{*}$-invariant topology. Then there is a natural $\mathbb{Z}=\Omega S^{1}$-linear structure on $(\gamma_{\mathbb{C}})_{*}\mu sh_{W\times\{0\}}$ and an equivalence
(22) $(\gamma_{\mathbb{C}})_{*}\left(\mu sh_{W\times\{0\}}\otimes_{\Omega S^{1}}\bullet\right)\simeq(\gamma_{\mathbb{C}})_{*}\mu sh_{W}.$
Furthermore, this equivalence respects complex constructibility.
Proof.
Up to replacing $W$ by a $\mathbb{C}^{*}$-invariant open, it is enough to exhibit the $\Omega S^{1}$-linear structure on $\mu sh_{W\times\{0\}}(W\times\{0\})$, and to prove the equivalence (22) on global sections.
By our assumption on Liouville vector field, the change of variables (21) is global:
(23) $(\widetilde{W\times\mathbb{C}}=W\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz))$ $\xrightarrow{\simeq}(W\times\mathbb{C}\times\mathbb{C}^{*},\lambda+wdz).$
(24) $\mu sh_{W\times 0\times\mathbb{C}^{*}}=\mu sh_{W}\boxtimes\mu sh_{\mathbb{C}^{*}\subset T^{*}\mathbb{C}^{*}}.$
By definition
[…]
where $\mu sh_{\pi^{-1}(W\times\{0\})}$ denotes the sheaf of full subcategories of $\mu sh_{\widetilde{W\times\mathbb{C}}}$ on objects whose support is contained in $\pi^{-1}(W\times\{0\})$; see (19). But by (23) and (24), we have
$\left(\pi_{*}\mu sh_{\pi^{-1}(W\times\{0\})}\right)(\widetilde{W\times\mathbb{C}})=\mu sh_{W}(W)\otimes loc(\mathbb{C}^{*}).$
@@
Let $\mathbb{C}^{*}\times W\to W,(\theta,z)\mapsto\theta\cdot z$ be the weight-11 $\mathbb{C}^{*}$ action on $W$. Then $(\theta;x,z)\mapsto(\theta\cdot x,\theta z)$ defines a $\mathbb{C}^{*}$ action on $W\times\mathbb{C}$ by contactomorphism, which fixes $W\times\{0\}$ set-wise. (22) amounts to taking invariants of this action; see [33, Sec. 6].